# LLM-wrapper: Black-Box Semantic-Aware Adaptation of Vision-Language Models for Referring Expression Comprehension

### v2

@@ Abstract
Vision Language Models (VLMs) have demonstrated remarkable capabilities in various open-vocabulary tasks, yet their zero-shot performance lags behind task-specific fine-tuned models, particularly in complex tasks like Referring Expression Comprehension (REC). Fine-tuning usually requires ‘white-box’ access to the model’s architecture and weights, which is not always feasible due to proprietary or privacy concerns. In this work, we propose LLM-wrapper, a method for ‘black-box’ adaptation of VLMs for the REC task using Large Language Models (LLMs). LLM-wrapper capitalizes on the reasoning abilities of LLMs, improved with a light fine-tuning, to select the most relevant bounding box matching the referring expression, from candidates generated by a zero-shot black-box VLM. Our approach offers several advantages: it enables the adaptation of closed-source models without needing access to their internal workings, it is versatile as it works with any VLM, it transfers to new VLMs, and it allows for the adaptation of an ensemble of VLMs. We evaluate LLM-wrapper on multiple datasets using different VLMs and LLMs, demonstrating significant performance improvements and highlighting the versatility of our method. While LLM-wrapper is not meant to directly compete with standard white-box fine-tuning, it offers a practical and effective alternative for black-box VLM adaptation. The code will be open-sourced.
@@ 1 Introduction
Vision Language Models (VLMs), a class of foundation models (Bommasani et al., 2021), trained on large-scale and diverse tasks and datasets, have shown remarkable abilities to solve various open-vocabulary tasks, as image captioning (Xiao et al., 2023; Li et al., 2023), visual question answering (Alayrac et al., 2022; Liu et al., 2023; Li et al., 2023), text-image retrieval (Radford et al., 2021; Zhai et al., 2023; Li et al., 2023), object detection (Liu et al., 2024c; Xiao et al., 2023; Cheng et al., 2024), or semantic segmentation (Ding et al., 2023; Xiao et al., 2023). Recent VLMs show promising zero-shot generalization abilities to new tasks and data domains (Wei et al., 2022; Alayrac et al., 2022). However, there is still a significant performance gap between zero-shot VLMs and those that have been specifically trained or adapted for a particular task and data domain. This work focuses on the challenging open-vocabulary detection task of Referring Expression Comprehension (REC) (Mao et al., 2016), which involves localizing an object in an image based on a complex textual query, requiring both spatial and semantic reasoning. While zero-shot VLMs can typically detect most objects mentioned in the query with reasonably accurate bounding boxes and labels, they struggle to identify only the described object. Moreover, VLMs often have difficulty understanding complex descriptions that involve relations between objects, attributes, or negations (Xie et al., 2023; Yao et al., 2024).
To improve performance, VLMs are typically fine-tuned on the specific REC task and corresponding datasets. This fine-tuning is usually done in a ‘white-box’ manner, with full access to the model’s architecture and weights for back-propagation. However, this process requires expertise to design fine-tuning objectives and optimize hyper-parameters, specific to each VLM and downstream task. Moreover, white-box fine-tuning is not always feasible. Some models are closed-source, either because they are proprietary and released behind APIs (Ren et al., 2024), or because they are trained on private data, making their weights and gradients inaccessible. While companies may provide APIs for adapting proprietary models, e.g., (OpenAI, 2024), these solutions are limited to predefined scopes and require sharing data with external private companies, raising legal and privacy concerns.
@@
LLM-wrapper offers several advantages, summarized in Table 1. First, since the adaptation is done in a black-box manner, it allows for the adaptation of closed-source models, without needing access to the model’s architecture, weights, and gradients. This makes black-box fine-tuning versatile and easy to use, as it can be applied to any model without requiring specific model knowledge or assumptions about the model’s architecture. Additionally, as the adaptation is delegated to the LLM, which learns the specifics of the task, the adaptation retains the pre-trained knowledge of the original VLM, and we find that LLM-wrapper transfers well to other VLMs and generalizes to model updates. Finally, LLM-wrapper enables the adaptation of an ensemble of VLMs, thereby allowing to combine strengths from multiple models.
We evaluate LLM-wrapper on three datasets – RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), RefCOCOg (Mao et al., 2016) – using two notably different VLMs – Grounding-DINO (Liu et al., 2024c) and Florence-2 (Xiao et al., 2023). We also experiment with two LLMs for the adaptation: the recent Mixtral 8x7B (Jiang et al., 2024) and Llama 3 8B (Dubey et al., 2024). While LLM-wrapper is not meant to outperform standard white-box fine-tuning, we show that LLM-wrapper significantly enhances the VLM’s performances, across all combinations of VLMs, LLMs and datasets, thus demonstrating the versatility of our method. Notably, on RefCOCOg, our most challenging dataset with respect to semantic understanding, LLM-wrapper improves the results of zero-shot VLMs with gains ranging from +9.5 up to +18.0 P@1.
@@
Fine-tuning generalization Cannot directly transfer to nor ensemble other models’ outputs Transfers well from one VLM to another and can ensemble outputs
@@ 2 Related work
Referring Expression Comprehension (REC) (Kazemzadeh et al., 2014; Mao et al., 2016; Qiao et al., 2021) is the task of identifying objects in an image based on referring expressions. Typically, REC involves selecting the best region from a set of region proposals extracted from the image, guided by a referring expression. This task is challenging because the referring expression can range from a short phrase (Kazemzadeh et al., 2014; Yu et al., 2016) to a long sentence that may require multi-step reasoning (Mao et al., 2016). REC is distinct from similar tasks such as visual grounding (Rohrbach et al., 2016), where multiple object regions described by multiple noun phrases must be localized in an image. It also differs from object detection (Girshick, 2015), which uses predefined categories instead of natural language expressions. The recent surge of interest in VLMs, starting from CLIP (Radford et al., 2021) to the recent Florence-2 (Xiao et al., 2023) has significantly improved performance on tasks requiring both vision and language. For REC, several approaches have used VLMs. For instance, Grounding-DINO (Liu et al., 2024c) relies on multiple stages of modality fusion to align visual and textual features. The more recent Florence-2 (Xiao et al., 2023) is a sequence-to-sequence model trained on a huge collection of data. However, despite being trained on extensive amounts of image-text data (e.g., 126 million images with 500 million to 3.6 billion annotations for Florence-2), these models’ zero-shot performances on REC are sub-optimal compared to their performances when including REC data in their training set or when fine-tuned (see Section 4.2). Our approach, LLM-wrapper, addresses this clear limitation in a black-box manner, without the need to re-train the VLM with task-specific data.
@@
While retraining the entire VLM on a new task or data domain is computationally expensive (Liu et al., 2023; Wang et al., 2023; Chen et al., 2023; Xiao et al., 2023), fine-tuning offers a more efficient alternative. However, even fine-tuning can be costly. To address this, parameter-efficient fine-tuning, e.g., LoRa (Hu et al., 2022), DoRa (Liu et al., 2024a), or VeRa (Kopiczko et al., 2024), and soft prompt (Li & Liang, 2021) learning have been proposed. These methods avoid updating the model’s pre-trained weights but require access to the model’s architecture and gradients.
A few recent methods propose strategies to adapt VLMs when back-propagation through the VLM is not feasible (Ouali et al., 2023; Yu et al., 2023; Liu et al., 2024b; Oh et al., 2023). Both Ouali et al. (2023) and Yu et al. (2023) focus on CLIP-based methods and adapt the VLM by learning either a feature projection or soft prompts. They however require access to the inner representations, which is generally not possible with APIs. Liu et al. (2024b) optimize the input prompt template, but also focuses on CLIP-based models. Instead, we target VLMs that perform open-vocabulary detection, e.g., Grounding-DINO and Florence-2. Finally, (Oh et al., 2023) adapt a VLM in a black-box manner by modifying the image input. However, this method is designed to address visual domain gaps and is not directly applicable to adapt a VLM to a new task. Moreover, it requires multiple API calls for each training image. In this work, we propose a strategy to adapt VLMs, and more specifically open-vocabulary detectors, to a new task, in a complete black-box fashion. Our method does not require access to the model’s intermediate representations or gradients, runs a single API call for each (text-image) input pair, and enables adaptation to a new task (REC in our case) that the VLM was not originally trained for.
Figure 1: Illustration of LLM-wrapper. Our method adapts a black-box VLM for the REC task. The output of the VLM is translated into a natural language text that is used to prompt an LLM. The LLM is then tasked with identifying the box that best matches the referring expression among the given candidates. The LLM must learn to identify the subject of the query and to disambiguate the correct object from other distractor objects (e.g., several plates with ‘flower details’).
@@
In this section, we present LLM-wrapper, a novel LLM-based approach to adapt an open-ended VLM for the REC task. We present the general idea in Section 3.1, the construction of the prompt for the LLM in Section 3.2, the fine-tuning details in Section 3.3. Lastly, we present the benefits of having a language-based adaptation of VLMs’ outputs in Section 3.4.
@@ 3.2 Prompt construction
The key idea of our method is to convert the VLM’s outputs into natural language. To achieve this, we list all predicted outputs in the LLM prompt, including their box coordinates, labels, and, when applicable, prediction scores (displayed below in light gray). This allows the ‘blind’ LLM, which only reads text, to understand the scene and reason about the image. The prompt then reminds the query and asks the LLM to select the best matching box.
For instance, given the query (in green in Figure 1), and the associated outputs (e.g., ‘flower’, ‘plate with flower details’, ‘yellow bowl’, etc.), we ask the LLM for the best matching box index, as follows:
@@ 3.3 Fine-tuning the LLM
3.4 Benefits of wrapping VLMs’ outputs
Our approach aims at adapting VLMs in a black-box manner. We convert their outputs into natural language and wrap them with a language model, which has several benefits. First, by working on the outputs of the VLM rather than its inputs, we can back-propagate into the LLM. Second, during fine-tuning, LLM-wrapper learns general knowledge about the task at hand. This allows us to transfer LLM-wrapper from one VLM to another. For example, we can leverage the updated versions of models available via API, such as moving from Grounding-DINO to Grounding-DINO-1.5, without retraining the language model (see Section 4.3). Finally, since the input format of LLM-wrapper is free text, the method can easily adapt to a varying number of input bounding boxes. For instance, our method can naturally extend to an ensemble of multiple VLMs (see Section 4.3). We observe that this ‘ensembling’ approach is particularly useful to combine the VLMs’ strengths and mitigate their different biases. Overall, the input format of our method makes it flexible for new use cases.
@@ 4.1.1 Referring Expression Comprehension (REC) task
Datasets. We evaluate LLM-wrapper on the REC task. In REC, given an input pair (image, query), a model is expected to predict a single bounding box around the object described in the query, as illustrated in Figure 1. We use three standard datasets for REC, RefCOCO (Yu et al., 2016), RefCOCO+ (Yu et al., 2016), and RefCOCOg (Mao et al., 2016) which is the most challenging as it contains longer and more complex descriptions (8.4 words per query in average). All of these datasets contain multiple distractor objects in each image for the given text query. Dataset’s statistics are reported in Table 2.
Table 2: REC datasets statistics.
@@ 4.1.2 The VLMs
We evaluate the impact of LLM-wrapper on two different VLMs. In all cases, we use the official model checkpoints and unless specified otherwise, we use the model versions that are not fine-tuned for the REC task, meaning that the models have not been exposed to any of the RefCOCO/+/g datasets. Using the model setup described below, we feed from 2 to 20 boxes to LLM-wrapper per prompt.
Grounding-DINO (GD) (Liu et al., 2024c) aligns visual queries with text through stages of modality fusion and contrastive learning. Initially designed for open-vocabulary detection tasks, the model produces 900 bounding boxes, achieving a high recall but an under-performing precision on the REC task. We propose a straightforward improvement for REC, Grounding-DINO-REC (GDrec), in which instead of assigning to each box the maximum similarity score with respect to any query word as confidence score, we assign the similarity score corresponding to the query’s subject [^1] The subject is defined as the first noun group identified by Spacy’s dependency parser (Honnibal et al., 2020).. GDrec significantly outperforms GD on all datasets, and particularly on RefCOCOg, where the noun identification is more challenging, with e.g., +7 P@1 for zero-shot models as shown in Table 3. To ensure a rather short prompt, we limit the number of box proposals by setting the box confidence score threshold to $0.15$ for GDrec and to $0.2$ for GD. GD / GDrec, not fine-tuned, use a SwinT(T) backbone while the fine-tuned versions are based on SwinB(B), as they are the only publicly available models. We also run experiments on a subset of RefCOCOg using Grounding-DINO 1.5 (Ren et al., 2024) (GD-1.5), a recent detector behind API (online at: Grounding-DINO-1.5-API), which provides 300 free API calls. GD-1.5 extends GD with a larger backbone, ViT-L (Fang et al., 2024), and training dataset.
Florence-2 (Flo2) (Xiao et al., 2023) is a sequence-to-sequence multi-task model. We use the Florence-2 Large version from the Hugging Face Hub. It is composed of a DaViT vision encoder (Ding et al., 2022) and a multi-modal encoder-decoder. Flo2 can be prompted for several tasks, and we keep and concatenate the boxes from the ‘open vocabulary detection’ and the ‘phrase grounding’ task modes.

### v3

@@ Abstract
Vision Language Models (VLMs) have demonstrated remarkable capabilities in various open-vocabulary tasks, yet their zero-shot performance lags behind task-specific fine-tuned models, particularly in complex tasks like Referring Expression Comprehension (REC). Fine-tuning usually requires ‘white-box’ access to the model’s architecture and weights, which is not always feasible due to proprietary or privacy concerns. In this work, we propose LLM-wrapper, a method for ‘black-box’ adaptation of VLMs for the REC task using Large Language Models (LLMs). LLM-wrapper capitalizes on the reasoning abilities of LLMs, improved with a light fine-tuning, to select the most relevant bounding box matching the referring expression, from candidates generated by a zero-shot black-box VLM. Our approach offers several advantages: it enables the adaptation of closed-source models without needing access to their internal workings, it is versatile as it works with any VLM, it transfers to new VLMs and datasets, and it allows for the adaptation of an ensemble of VLMs. We evaluate LLM-wrapper on multiple datasets using different VLMs and LLMs, demonstrating significant performance improvements and highlighting the versatility of our method. While LLM-wrapper is not meant to directly compete with standard white-box fine-tuning, it offers a practical and effective alternative for black-box VLM adaptation. The code and the checkpoints are available at https://github.com/valeoai/LLM_wrapper.
@@ 1 Introduction
Vision Language Models (VLMs), a class of foundation models (Bommasani et al., 2021), trained on large-scale and diverse tasks and datasets, have shown remarkable abilities to solve various open-vocabulary tasks, as image captioning (Xiao et al., 2024; Li et al., 2023), visual question answering (Alayrac et al., 2022; Liu et al., 2023; Li et al., 2023), text-image retrieval (Radford et al., 2021; Zhai et al., 2023; Li et al., 2023), object detection (Liu et al., 2024c; Xiao et al., 2024; Cheng et al., 2024), or semantic segmentation (Ding et al., 2023; Xiao et al., 2024). Recent VLMs show promising zero-shot generalization abilities to new tasks and data domains (Wei et al., 2022; Alayrac et al., 2022). However, there is still a significant performance gap between zero-shot VLMs and those that have been specifically trained or adapted for a particular task and data domain. This work focuses on the challenging open-vocabulary detection task of Referring Expression Comprehension (REC) (Mao et al., 2016), which involves localizing an object in an image based on a complex textual query, requiring both spatial and semantic reasoning. While zero-shot VLMs can typically detect most objects mentioned in the query with reasonably accurate bounding boxes and labels, they struggle to identify only the described object. Moreover, VLMs often have difficulty understanding complex descriptions that involve relations between objects, attributes, or negations (Xie et al., 2023; Yao et al., 2024).
To improve performance, VLMs are typically fine-tuned on the specific REC task and corresponding datasets. This fine-tuning is usually done in a ‘white-box’ manner, with full access to the model’s architecture and weights for back-propagation. However, this process requires expertise to design fine-tuning objectives and optimize hyper-parameters, specific to each VLM and downstream task. Moreover, white-box fine-tuning is not always feasible. Some models are closed-source, either because they are proprietary and released behind APIs (Ren et al., 2024a), or because they are trained on private data, making their weights and gradients inaccessible. While companies may provide APIs for adapting proprietary models, e.g., (OpenAI, 2024), these solutions are limited to predefined scopes and require sharing data with external private companies, raising legal and privacy concerns.
@@
LLM-wrapper offers several advantages, summarized in Table 1. First, since the adaptation is done in a black-box manner, it removes the need for back-propagation through the VLM and allows for the adaptation of closed-source models, without requiring access to the model’s architecture, weights, or gradients. This makes black-box fine-tuning versatile and easy to use, as it can be applied to any model without requiring specific model knowledge or assumptions about the model’s architecture. Additionally, as the adaptation is delegated to the LLM via text, LLM-wrapper retains the pre-trained knowledge of the original VLM. We find that LLM-wrapper transfers well to other VLMs, generalizes to model updates, and can adapt effectively to new datasets without additional fine-tuning. Finally, LLM-wrapper enables the adaptation of an ensemble of VLMs, leveraging the flexibility of text-based adaptation to handle varying numbers of bounding boxes, thereby combining strengths from multiple models.
We experiment with LLM-wrapper on three classic REC datasets — RefCOCO, RefCOCO+ (Kazemzadeh et al., 2014), RefCOCOg (Mao et al., 2016) — and on Talk2Car (Deruyttere et al., 2019), using two notably different VLMs — Grounding-DINO (Liu et al., 2024c) and Florence-2 (Xiao et al., 2024). Additionaly, we evaluate LLM-wrapper on the recent and challenging HC-RefLoCo (Wei et al., 2024) benchmark. We also experiment with two LLMs for the adaptation: Mixtral 8x7B (Jiang et al., 2024) and Llama 3 8B (Dubey et al., 2024). While LLM-wrapper is not meant to outperform standard white-box fine-tuning, we show that LLM-wrapper significantly enhances the VLM’s performances, across all combinations of VLMs, LLMs and datasets, thus demonstrating the versatility of our method. Notably, on RefCOCOg and Talk2Car, our most challenging benchmarks with respect to semantic understanding, LLM-wrapper improves the results of zero-shot VLMs with gains ranging from +8.7 up to +18.3 P@1.
@@
Fine-tuning generalization Limited: no ensembling, no transfer to new VLMs or datasets Flexible: supports ensembling and transfers across VLMs and datasets
@@ 2 Related work
Referring Expression Comprehension (REC) (Kazemzadeh et al., 2014; Mao et al., 2016; Qiao et al., 2021) is the task of identifying objects in an image based on referring expressions. Typically, REC involves selecting the best region from a set of region proposals extracted from the image, guided by a referring expression. This task is challenging because the referring expression can range from a short phrase (Kazemzadeh et al., 2014) to a long sentence that may require multi-step reasoning (Mao et al., 2016). REC is distinct from similar tasks such as visual grounding (Rohrbach et al., 2016), where multiple object regions described by multiple noun phrases must be localized in an image. It also differs from object detection (Girshick, 2015), which uses predefined categories instead of natural language expressions. The recent surge of interest in VLMs, starting from CLIP (Radford et al., 2021) to the recent Florence-2 (Xiao et al., 2024) has significantly improved performance on tasks requiring both vision and language. For REC, several approaches have used VLMs. For instance, Grounding-DINO (Liu et al., 2024c) relies on multiple stages of modality fusion to align visual and textual features, while Florence-2 (Xiao et al., 2024) is a sequence-to-sequence model trained on a huge collection of data. However, despite being trained on extensive amounts of image-text data (e.g., 126 million images with 500 million to 3.6 billion annotations for Florence-2), these models’ zero-shot performances on REC are sub-optimal compared to their performances when including REC data in their training set or when fine-tuned (see Section 4.2). Our approach, LLM-wrapper, addresses this clear limitation in a black-box manner, without the need to re-train the VLM with task-specific data.
@@
While retraining the entire VLM on a new task or data domain is computationally expensive (Liu et al., 2023; Wang et al., 2023; Chen et al., 2023; Xiao et al., 2024), fine-tuning offers a more efficient alternative. However, even fine-tuning can be costly. To address this, parameter-efficient fine-tuning, e.g., LoRa (Hu et al., 2022), DoRa (Liu et al., 2024a), or VeRa (Kopiczko et al., 2024), and soft prompt (Li & Liang, 2021) learning have been proposed. These methods avoid updating the model’s pre-trained weights but require access to the model’s architecture and gradients.
A few recent methods propose strategies to adapt VLMs when back-propagation through the VLM is not feasible (Ouali et al., 2023; Yu et al., 2023; Liu et al., 2024b; Oh et al., 2023; Wan et al., 2024). Both Ouali et al. (2023) and Yu et al. (2023) focus on CLIP-based methods and adapt the VLM by learning either a feature projection or soft prompts. They however require access to the inner representations, which is generally not possible with APIs. Liu et al. (2024b) optimize the input prompt template, but also focuses on CLIP-based models. Instead, we target all VLMs that perform open-vocabulary detection, e.g., Grounding-DINO and Florence-2. Oh et al. (2023) adapt a VLM in a black-box manner by modifying the image input. However, this method is designed to address visual domain gaps and is not directly applicable to adapt a VLM to a new task. Moreover, it requires multiple API calls for each training image. Finally, Wan et al. (2024) introduce an adaptation method for frozen VLMs, applicable to REC. For each candidate bounding box, they perform a VLM inference to compute the probability of the input text based on a modified image where the box has been blacked out. The candidate that achieves the highest probability contrast with respect to the unmodified input image is chosen. Though training-free, this approach requires the access to the model’s output probability distribution and performs an additional VLM inference for each candidate bounding box. In this work, we propose a strategy to adapt VLMs, and more specifically open-vocabulary detectors, to a new task, in a complete black-box fashion. Our method does not require access to the model’s intermediate representations, output distribution, or gradients, runs a single inference/API call for each (text-image) input pair, and enables adaptation to a new task (REC in our case) that the VLM was not originally trained for.
Figure 1: Illustration of LLM-wrapper. Our method adapts a black-box VLM for the REC task. The VLM output is translated into natural language to prompt an LLM. The latter is tasked with identifying the box that best matches the query among the given candidates. The LLM must learn to identify the query’s subject and to disambiguate the correct object from distractors (e.g., several plates with ‘flower details’).
@@
We present LLM-wrapper, a novel LLM-based approach to adapt VLMs for the REC task. We detail the general idea in Section 3.1, the prompt construction in Section 3.2 and the LLM fine-tuning in Section 3.3.
@@ 3.2 Prompt construction
The key idea of our method is to convert the VLM’s outputs into natural language. To achieve this, we list all predicted outputs in the LLM prompt, including their box coordinates, labels, and, when applicable, prediction scores (displayed below in light gray). This allows the ‘blind’ LLM, which only reads text, to understand the scene and reason about the image. The prompt then reminds the query and asks the LLM to select the best matching box. For instance, given the query (in green in Figure 1), and the associated outputs (e.g., ‘flower’, ‘plate with flower details’, ‘yellow bowl’, etc.), we ask the LLM for the best matching box index, as follows:
@@ 4.1.1 Referring Expression Comprehension (REC) task
Datasets. We evaluate LLM-wrapper on the REC task. In REC, given an input pair (image, query), a model is expected to predict a single bounding box around the object described in the query, as illustrated in Figure 1. We use three standard datasets for REC: RefCOCO, RefCOCO+ (Kazemzadeh et al., 2014), and RefCOCOg (Mao et al., 2016) which is more challenging as it contains longer descriptions (8.3 words in average). We also use the driving, non human-centric, Talk2Car (Deruyttere et al., 2019) REC dataset. Additionally, we evaluate LLM-wrapper on HC-RefLoCo (Wei et al., 2024) for zero-shot dataset transfer experiments. All of these datasets contain multiple distractor objects for the given text query. Dataset statistics are given in Table 2.
Table 2: Dataset statistics.
@@ 4.1.2 The VLMs
We evaluate the impact of LLM-wrapper on two different VLMs. In all cases, we use the official model checkpoints and unless specified otherwise, we use the model versions that are not fine-tuned for the REC task, meaning that the models have not been exposed to any of the RefCOCO/+/g and Talk2Car datasets. Using the model setup described below, we feed from 2 to 45 boxes to LLM-wrapper per prompt.
Grounding-DINO (GD) (Liu et al., 2024c) aligns visual queries with text through stages of modality fusion and contrastive learning. Initially designed for open-vocabulary grounding, the model produces 900 bounding boxes, achieving a high recall but an under-performing precision on the REC task. For REC, we use GD by selecting the bounding box with the highest score relative to any token in the query. However, this method sometimes selects boxes based on query parts unrelated to the main object, such as other nouns in the sentence. To address this, we introduce Grounding-DINO-REC (GDrec), a more targeted approach for REC. GDrec identifies the query’s subject, defined as the first noun group detected by SpaCy’s dependency parser (Honnibal et al., 2020), and selects the bounding box that scores best against it. GDrec outperforms GD on RefCOCO/+/g datasets, and particularly on RefCOCOg, where the noun identification is more challenging, with e.g., +9.1 P@1 (test) for zero-shot models as shown in Table 3. To ensure a rather short prompt, we limit the number of box proposals by setting the box confidence score threshold to $0.15$ for GDrec and to $0.2$ for GD. GD / GDrec, not fine-tuned, use a SwinT(T) backbone while the fine-tuned versions are based on SwinB(B), as they are the only publicly available models. We also run experiments on a subset of RefCOCOg using Grounding-DINO 1.5 (Ren et al., 2024a) (GD-1.5), a recent detector behind API (online at: Grounding-DINO-1.5-API), which provides 300 free API calls. GD-1.5 extends GD with a larger backbone, ViT-L (Fang et al., 2024), and training dataset.
Florence-2 (Flo2) (Xiao et al., 2024) is a sequence-to-sequence multi-task model. We use the Florence-2 Large version from Hugging Face. It is composed of a DaViT vision encoder (Ding et al., 2022) and a multi-modal encoder-decoder. Flo2 can be prompted to deal with different grounding tasks. To evaluate Flo2 on the REC task without adaptation, we use the box predicted for the ‘open vocabulary detection’ task. For LLM-wrapper’s adaptation of Flo2, we keep and concatenate the boxes predicted for both the ‘open vocabulary detection’ and the ‘phrase grounding’ task modes.